Calculate the critical micelle concentration for a methane molecule (r=0.2nm) in water (interface energy with water = 50mJ/m2) at 350K. (4 points)
Answer #1

Critical Micelle Concentration (CMC) is the concentration of surfactants above which micelles are spontaneously formed.

We know that,

Given that,

Temperature, T = 350 K

The interface energy. G = 50mJ/m2

Radius of methane molecule,r = 0.2nm

We have the eqaution,

G * Sc = kT ln CMC, where k is boltzmann constant

i.e, 50 * (0.2)2 = 1.38 * 10-23 * 350 * ln CMC

i.e, ln CMC = 0.00414 * 1023